OpenStudy (anonymous):

how do simplify this further: (2x-7/x-2) - (x-5/x-2) = (2x-7/x-2) + (x+5/x-2)?

OpenStudy (mertsj):

\[\frac{2x-7}{x-2}-\frac{x-5}{x-2}=\frac{2x-7-x+5}{x-2}=\frac{x-2}{x-2}=1\]

OpenStudy (anonymous):

why is that the numerator? i got \[2x ^{2}-2\]

OpenStudy (mertsj):

Consider only the numerators: 2x-7-(x-5)=2x-7-x+5=x-2

OpenStudy (anonymous):

so its basically 2+(-1)=1 so it comes out to be x-2 ?

OpenStudy (mertsj):

yes. 2x-x=x

OpenStudy (mertsj):

Also -7+5=-2

OpenStudy (anonymous):

i understand the -2 part. thanks!

OpenStudy (anonymous):

((2x−7) - (x-2))/(x-2) = (2x - 7 - x -2)/(x-2) Remember that the denominator is the same so it can be subtracted or add the numerators (x+1)/(1-x) = (1/(1-x)) - ((x)/(1-x))

OpenStudy (anonymous):

Ugh I mean (1-x)/(1-x) = (1/(1-x)) - ((x)/(1-x))